# Masked Generative Policy for Robotic Control

### v1

@@ 1 Introduction
Enabling robots to perform complex manipulation tasks directly from high-dimensional sensory inputs, such as vision, remains a challenge in robotics and artificial intelligence. Imitation learning has emerged as a promising and data-efficient paradigm to tackle this challenge, bypassing the complex modeling process of reinforcement learning by directly leveraging human demonstrations (zare2024survey). Moving beyond simple state-action mapping, recent advancements in learning visuomotor policies have formulated the problem as training a generative model over action sequences conditioned on observations. These approaches typically use either (1) diffusion policies, which generate behavior via a conditional denoising process in robot action space (janner2022planning; chi2023diffusion; ze20243d), or (2) autoregressive policies, which treat actions as discrete tokens and model these tokens with a GPT-like transformer (mete2024quest).
State-of-the-art generative policies face inherent trade-offs that limit their use in closed-loop, real-time robotic control. Diffusion-based methods require multiple denoising steps per action, making them computationally slow. Consistency Policy (prasad2024consistency) and Flow Policy (zhang2025flowpolicy) have aimed to speed up sampling but they either require additional distillation or compromise sample quality. Autoregressive policies, on the other hand, sample one token per forward pass; therefore, latency scales with sequence length. Moreover, without memory, these policies lack robustness to missing observations and fail in non-Markovian tasks.
@@
Motivated by these limitations, our goal is to develop generative policies that simultaneously address the inference time bottlenecks of iterative sampling and the robustness challenges inherent in non-Markovian, long-horizon manipulation tasks. For this, we introduce Masked Generative Policy (MGP), which achieves low inference latency and high task success rates while supporting rapid plan edits during execution. MGP adapts masked generative transformers (chang2022maskgit) to model latent action representations. Actions are encoded as discrete token sequences via vector quantization and a conditional masked transformer reconstructs complete sequences from partial masks conditioned on observations.
@@ 2 Related work
Imitation learning enables an agent to acquire an expert policy without explicit rewards by training on expert (schaal1996learning; ho2016generative). This is carried out typically via behavior cloning (NIPS1988_812b4ba2; 10.5555/3304652.3304697; florence2021implicit) or inverse reinforcement learning (10.5555/645529.657801; pengvariational; fu2018learning). With the advent of deep learning, directly learning expert behaviors through time-series models has gradually become the dominant paradigm (zhao2023learning; florence2019self).
Recent works (chi2023diffusion; xian2023chaineddiffuser) have shown that diffusion models (10.5555/3495724.3496298) can learn visuomotor policies by treating action synthesis as a conditional denoising process. Similarly, ze20243d and ke20243d have proposed to enhance the performance of diffusion-based policies by incorporating multimodal information such as point clouds and text instructions. Meanwhile, diffusion models have also been widely applied to trajectory planning (zheng2025diffusion; xiao2023safediffuser). Due to the long trajectory of denoising, several studies have attempted to improve its sampling efficiency, for example, FlowPolicy (zhang2025flowpolicy) and Consistency Policy (prasad2024consistency). However, these methods either compromise sampling quality or require additional distillation.
@@
The stepwise prediction of actions in autoregressive models aligns naturally with the sequential operation of robotic task execution (zhang2025autoregressive). Therefore, researchers have explored a variety of autoregressive models for imitation learning (sun2017deeply; xie2020deep), with transformer-based architectures gradually emerging as the dominant approach due to their strong sequence modeling capacity (cui2023from; Brohan-RSS-23; chebotar2023qtransformer; zhao2023learning). More recently, large language models have extended transformer-based architectures and introduced the next-token prediction paradigm into this field, where GPT-like pipelines encode actions into discrete embeddings via a tokenizer, and the transformer models the correspondence between these embeddings and observations (shafiullah2022behavior). PRISE (zheng2024prise),VQ-BeT (lee2024behavior) and QueST (mete2024quest) have achieved improved task success rates and generalization by employing novel discrete representations. Other work like Chain-of-Action (pan2024chain) further extends autoregressive approaches by decoding trajectories in reverse from a goal keyframe to mitigate compounding error. However, autoregressive policies have inherent structural limitations: (i) next-token prediction leads to multi-step iterations in long-horizon action generation, and (ii) immutability of prefixes—any edit needs regenerating all subsequent tokens.
@@
The masking mechanism was initially employed as a regularization technique in deep learning (vincent2008extracting; pathak2016context; bao2022beit; he2022masked). Building on this foundation, MaskGIT (chang2022maskgit) and MUSE (chang2023muse) have established masked token modeling as a scalable, high-fidelity paradigm for image and text-to-image synthesis, while StyleDrop (sohn2023styledrop) demonstrates controllable, personalized generation via token-level conditioning. This paradigm has since extended beyond images. For example, MMM (pinyoanuntapong2024mmm) and MoMask (guo2024momask) have applied masked generative modeling to long-horizon human motion.
@@ 3 Methods
MGP is trained by imitation learning on a dataset of expert demonstrations. We first train an Action Tokenizer that learns a discrete representation of robot action sequences (Sec. 3.1). A Masked Generative Transformer (MGT) then learns to reconstruct masked sequences of action tokens (Sec. 3.2), allowing the generation of future actions conditioned on observations. Building upon MGT chang2022maskgit, we introduce two sampling paradigms, MGP-Short for Markovian (Sec. 3.3) and MGP-Long for Non-Markovian tasks (Sec. 3.4).
@@ 3.1 Action Tokenizer
We use a VQ-VAE (van2017neural) to obtain a discrete representation of actions. Specifically, the VQ-VAE takes a sequence of continuous-valued actions $\mathbf{a}$ as input and maps this to a shorter sequence of discrete action tokens, which can be reconstructed to the corresponding actions (Fig. 2-1). This creates a discrete latent space for the MGT to operate over.
@@
The training objectives of the model follow the VQ-VAE (van2017neural), which includes reconstruction loss and commitment loss:
@@
where $\rm sg$$\cdot$] is the stop-gradient operator, $\lambda_{\text{rec}}$ is $1$ and $\beta$ is $0.02$. We use exponential moving averages (EMA) for codebook updates, and resetting of inactive codewords to guarantee codebook usage (chang2022maskgit). Once trained, the weights of the VQ-VAE are frozen. During the training of the subsequent MGT, the VQ-VAE is employed to encode the actions from the training data and to decode the tokens generated by the MGT. In the inference phase, only the VQ-decoder is used to reconstruct actions from the tokens sampled by the MGT.
@@ 3.2 Masked Generative Transformer
MGT is required to generate $N$ future action tokens sequence $\mathbf{y}^{0:N}$$y_{n}]^{N}_{n=0}$ conditioned on the observed inputs $O_{t}$ and the robot’s historical states $s_{t}$ from unknown tokens by a few gradual refinement steps. Initially, a special learnable token $\left[\mathrm{MASK}\right]$ is introduced to represent unknown tokens. In addition, [END] and [PAD] tokens are used to indicate the termination and padding of a token sequence, respectively (pinyoanuntapong2024mmm). The pipeline is illustrated in Fig. 2-2.
@@ 3.3 Short horizon Mask-and-Refine Sampling (MGP-Short)
MGP-Short only samples the tokens conditioned on the current observation $c_{t}$ obtained from the perception encoder. This procedure involves 2 iterations. For the first iteration, at time step $t$, the masked token sequence, together with $c_{t}$, is fed into the transformer to generate a first estimate of the token logits in parallel. The $n$-th token index is calculated from the raw logits $e_{n}$ by the Gumbel-Max trick  (huijben2022review):
@@ 4 Experiments and Results
In this section, we first evaluate on three standard benchmarks, Meta-World (yu2020meta), LIBERO-90 (liu2023libero), and LIBERO-Long, covering both short- and long-duration tasks under single-task and multi-task training (Section 4.1). Beyond these, we have designed three challenging evaluation environments that are important for long-horizon control and have repeatedly challenged prior methods: observation missing environments (Section 4.2); dynamic environments with moving objects/obstacles (Section 4.3); and two long-duration non-Markovian tasks (Section 4.4). Finally, we conduct four ablation studies (Section 4.5).
@@
We evaluate against ten baselines: diffusion methods-Diffusion Policy(chi2023diffusion), 3D Diffusion Policy (include simple DP3)(ze20243d)—two accelerated diffusion methods (Consistency Policy(prasad2024consistency) and FlowPolicy(zhang2025flowpolicy)) targeting faster sampling; and autoregressive methods—QueST(mete2024quest), VQ-BeT(lee2024behavior), PRISE(zheng2024prise), ACT(zhao2023learning), and ResNet-T(liu2023libero). Beyond the standard baselines, we also evaluate Full-Horizon DP3 (DP3-Full Seq.), a diffusion baseline that plans the entire action sequence from the initial observation and executes it open-loop (no replanning), similar to MGP-Full Seq.
@@ 4.1 Standard Benchmarks
We run single-task experiments on all 50 Meta-World tasks spanning Easy to Very Hard (seo2023masked). We use only short horizon methods on the Easy and Medium tasks, but long horizon on Hard and Very Hard tasks that benefit from a longer planning window. Ten expert demonstrations per task are generated using Meta-World heuristic policies. We evaluate 20 episodes every 1000 iterations using 3 seeds. We compute the average of top-5 success rates per task, per-step inference latency (ms) and per-sequence inference latency (ms).
@@
Our two-stage system is lightweight and fast: $\approx$7M parameters ($\approx$37× fewer than DP3 ze20243d’s $\approx$262M); training for 2000 epochs takes $\approx$55 minutes ($\approx$10 minutes for Stage one and $\approx$45 minutes for Stage two) vs. $\approx$3 hours on the same RTX 4090 setup. See Appendix Sec. C.6 for details.
@@ A.2 Standard Benchmarks
Meta-World (yu2020meta) is a simulated benchmark with a broad set of robotic manipulation tasks. We use the official difficulty classification that ranks tasks from easy to very hard (seo2023masked). Each demonstration contains 200 timesteps. In our experiments, we train 50 tasks on Meta-World separately.
We use LIBERO’s language-conditioned manipulation suites in simulation (liu2023libero). LIBERO-90 contains 90 single-goal, short-horizon tasks spanning diverse scenes (kitchen, living room, study) with rigid and articulated objects. Each task is accompanied by demonstrations ($\approx$50 per task) with multi-modal observations (RGB views and proprioception). LIBERO-LONG is a long-duration benchmark which comprises 10 long-horizon compositions, each formed by sequencing two LIBERO-90 goals, requiring multi-stage execution.
@@ C.8 Multimodality analysis
The Gumbel-Max trick ensures that token sampling still follows the logits’ softmax probability distribution, preserving diversity in the sampled results. To verify this, we conducted a qualitative evaluation of our model on the Push-T task florence2021implicit. In the initial stage, the control point, T-block, and target are placed symmetrically in fixed locations. MGP-Short then autoregressively generates 16 subsequent actions (4 tokens) and executes the first 8 actions based on the previous frame’s image and the control point position. The resulting trajectory plots (Fig. 9) show that MGP’s sampling still preserves the diversity distribution of the original data.

### v2

@@ 1 Introduction
Enabling robots to perform complex manipulation tasks directly from high-dimensional sensory inputs, such as vision, remains a challenge in robotics and artificial intelligence. Imitation learning has emerged as a promising and data-efficient paradigm to tackle this challenge, bypassing the complex modeling process of reinforcement learning by directly leveraging human demonstrations (Zare et al., 2024). Moving beyond simple state-action mapping, recent advancements in learning visuomotor policies have formulated the problem as training a generative model over action sequences conditioned on observations. These approaches typically use either (1) diffusion policies, which generate behavior via a conditional denoising process in robot action space (Janner et al., 2022; Chi et al., 2023; Ze et al., 2024), or (2) autoregressive policies, which treat actions as discrete tokens and model these tokens with a GPT-like transformer (Mete et al., 2024).
State-of-the-art generative policies face inherent trade-offs that limit their use in closed-loop, real-time robotic control. Diffusion-based methods require multiple denoising steps per action, making them computationally slow. Consistency Policy (Prasad et al., 2024) and Flow Policy (Zhang et al., 2025a) have aimed to speed up sampling but they either require additional distillation or compromise sample quality. Autoregressive policies, on the other hand, sample one token per forward pass; therefore, latency scales with sequence length. Moreover, without memory, these policies lack robustness to missing observations and fail in non-Markovian tasks.
@@
Motivated by these limitations, our goal is to develop generative policies that simultaneously address the inference time bottlenecks of iterative sampling and the robustness challenges inherent in non-Markovian, long-horizon manipulation tasks. For this, we introduce Masked Generative Policy (MGP), which achieves low inference latency and high task success rates while supporting rapid plan edits during execution. MGP adapts masked generative transformers (Chang et al., 2022) to model latent action representations. Actions are encoded as discrete token sequences via vector quantization and a conditional masked transformer reconstructs complete sequences from partial masks conditioned on observations.
@@ 2 Related work
Imitation learning enables an agent to acquire an expert policy without explicit rewards by training on expert (Schaal, 1996; Ho and Ermon, 2016). This is carried out typically via behavior cloning (Pomerleau, 1988; Torabi et al., 2018; Florence et al., 2021) or inverse reinforcement learning (Ng and Russell, 2000; Peng et al., 2018; Fu et al., 2018). With the advent of deep learning, directly learning expert behaviors through time-series models has gradually become the dominant paradigm (Zhao et al., 2023; Florence et al., 2019).
Recent works (Chi et al., 2023; Xian et al., 2023) have shown that diffusion models (Ho et al., 2020) can learn visuomotor policies by treating action synthesis as a conditional denoising process. Similarly, Ze et al. (2024) and Ke et al. (2024) have proposed to enhance the performance of diffusion-based policies by incorporating multimodal information such as point clouds and text instructions. Meanwhile, diffusion models have also been widely applied to trajectory planning (Zheng et al., 2025; Xiao et al., 2023). Due to the long trajectory of denoising, several studies have attempted to improve its sampling efficiency, for example, FlowPolicy (Zhang et al., 2025a) and Consistency Policy (Prasad et al., 2024). However, these methods either compromise sampling quality or require additional distillation.
@@
The stepwise prediction of actions in autoregressive models aligns naturally with the sequential operation of robotic task execution (Zhang et al., 2025b). Therefore, researchers have explored a variety of autoregressive models for imitation learning (Sun et al., 2017; Xie et al., 2020), with transformer-based architectures gradually emerging as the dominant approach due to their strong sequence modeling capacity (Cui et al., 2023; Brohan et al., 2023; Chebotar et al., 2023; Zhao et al., 2023). More recently, large language models have extended transformer-based architectures and introduced the next-token prediction paradigm into this field, where GPT-like pipelines encode actions into discrete embeddings via a tokenizer, and the transformer models the correspondence between these embeddings and observations (Shafiullah et al., 2022). PRISE (Zheng et al., 2024),VQ-BeT (Lee et al., 2024) and QueST (Mete et al., 2024) have achieved improved task success rates and generalization by employing novel discrete representations. Other work like Chain-of-Action (Pan et al., 2024) further extends autoregressive approaches by decoding trajectories in reverse from a goal keyframe to mitigate compounding error. However, autoregressive policies have inherent structural limitations: (i) next-token prediction leads to multi-step iterations in long-horizon action generation, and (ii) immutability of prefixes—any edit needs regenerating all subsequent tokens.
@@
The masking mechanism was initially employed as a regularization technique in deep learning (Vincent et al., 2008; Pathak et al., 2016; Bao et al., 2022; He et al., 2022). Building on this foundation, MaskGIT (Chang et al., 2022) and MUSE (Chang et al., 2023) have established masked token modeling as a scalable, high-fidelity paradigm for image and text-to-image synthesis, while StyleDrop (Sohn et al., 2023) demonstrates controllable, personalized generation via token-level conditioning. This paradigm has since extended beyond images. For example, MMM (Pinyoanuntapong et al., 2024) and MoMask (Guo et al., 2024) have applied masked generative modeling to long-horizon human motion.
@@ 3 Methods
MGP is trained by imitation learning on a dataset of expert demonstrations. We first train an Action Tokenizer that learns a discrete representation of robot action sequences (Sec. 3.1). A Masked Generative Transformer (MGT) then learns to reconstruct masked sequences of action tokens (Sec. 3.2), allowing the generation of future actions conditioned on observations. Building upon MGT Chang et al. (2022), we introduce two sampling paradigms, MGP-Short for Markovian (Sec. 3.3) and MGP-Long for Non-Markovian tasks (Sec. 3.4).
@@ 3.1 Action Tokenizer
We use a VQ-VAE (Van Den Oord et al., 2017) to obtain a discrete representation of actions. Specifically, the VQ-VAE takes a sequence of continuous-valued actions $\mathbf{a}$ as input and maps this to a shorter sequence of discrete action tokens, which can be reconstructed to the corresponding actions (Fig. 2-1). This creates a discrete latent space for the MGT to operate over.
@@
The training objectives of the model follow the VQ-VAE (Van Den Oord et al., 2017), which includes reconstruction loss and commitment loss:
@@
where $\rm sg$$\cdot$] is the stop-gradient operator, $\lambda_{\text{rec}}$ is $1$ and $\beta$ is $0.02$. We use exponential moving averages (EMA) for codebook updates, and resetting of inactive codewords to guarantee codebook usage (Chang et al., 2022). Once trained, the weights of the VQ-VAE are frozen. During the training of the subsequent MGT, the VQ-VAE is employed to encode the actions from the training data and to decode the tokens generated by the MGT. In the inference phase, only the VQ-decoder is used to reconstruct actions from the tokens sampled by the MGT.
@@ 3.2 Masked Generative Transformer
MGT is required to generate $N$ future action tokens sequence $\mathbf{y}^{0:N}$$y_{n}]^{N}_{n=0}$ conditioned on the observed inputs $O_{t}$ and the robot’s historical states $s_{t}$ from unknown tokens by a few gradual refinement steps. Initially, a special learnable token $\left[\mathrm{MASK}\right]$ is introduced to represent unknown tokens. In addition, [END] and [PAD] tokens are used to indicate the termination and padding of a token sequence, respectively (Pinyoanuntapong et al., 2024). The pipeline is illustrated in Fig. 2-2.
@@ 3.3 Short horizon Mask-and-Refine Sampling (MGP-Short)
MGP-Short only samples the tokens conditioned on the current observation $c_{t}$ obtained from the perception encoder. This procedure involves 2 iterations. For the first iteration, at time step $t$, the masked token sequence, together with $c_{t}$, is fed into the transformer to generate a first estimate of the token logits in parallel. The $n$-th token index is calculated from the raw logits $e_{n}$ by the Gumbel-Max trick  (Huijben et al., 2022):
@@ 4 Experiments and Results
In this section, we first evaluate on three standard benchmarks, Meta-World (Yu et al., 2020), LIBERO-90 (Liu et al., 2023), and LIBERO-Long, covering both short- and long-duration tasks under single-task and multi-task training (Section 4.1). Beyond these, we have designed three challenging evaluation environments that are important for long-horizon control and have repeatedly challenged prior methods: observation missing environments (Section 4.2); dynamic environments with moving objects/obstacles (Section 4.3); and two long-duration non-Markovian tasks (Section 4.4). Finally, we conduct four ablation studies (Section 4.5).
@@
We evaluate against ten baselines: diffusion methods-Diffusion Policy(Chi et al., 2023), 3D Diffusion Policy (include simple DP3)(Ze et al., 2024)—two accelerated diffusion methods (Consistency Policy(Prasad et al., 2024) and FlowPolicy(Zhang et al., 2025a)) targeting faster sampling; and autoregressive methods—QueST(Mete et al., 2024), VQ-BeT(Lee et al., 2024), PRISE(Zheng et al., 2024), ACT(Zhao et al., 2023), and ResNet-T(Liu et al., 2023). Beyond the standard baselines, we also evaluate Full-Horizon DP3 (DP3-Full Seq.), a diffusion baseline that plans the entire action sequence from the initial observation and executes it open-loop (no replanning), similar to MGP-Full Seq.
@@ 4.1 Standard Benchmarks
We run single-task experiments on all 50 Meta-World tasks spanning Easy to Very Hard (Seo et al., 2023). We use only short horizon methods on the Easy and Medium tasks, but long horizon on Hard and Very Hard tasks that benefit from a longer planning window. Ten expert demonstrations per task are generated using Meta-World heuristic policies. We evaluate 20 episodes every 1000 iterations using 3 seeds. We compute the average of top-5 success rates per task, per-step inference latency (ms) and per-sequence inference latency (ms).
@@
Our two-stage system is lightweight and fast: $\approx$7M parameters ($\approx$37× fewer than DP3 Ze et al. (2024)’s $\approx$262M); training for 2000 epochs takes $\approx$55 minutes ($\approx$10 minutes for Stage one and $\approx$45 minutes for Stage two) vs. $\approx$3 hours on the same RTX 4090 setup. See Appendix Sec. C.6 for details.
@@ A.2 Standard Benchmarks
Meta-World (Yu et al., 2020) is a simulated benchmark with a broad set of robotic manipulation tasks. We use the official difficulty classification that ranks tasks from easy to very hard (Seo et al., 2023). Each demonstration contains 200 timesteps. In our experiments, we train 50 tasks on Meta-World separately.
We use LIBERO’s language-conditioned manipulation suites in simulation (Liu et al., 2023). LIBERO-90 contains 90 single-goal, short-horizon tasks spanning diverse scenes (kitchen, living room, study) with rigid and articulated objects. Each task is accompanied by demonstrations ($\approx$50 per task) with multi-modal observations (RGB views and proprioception). LIBERO-LONG is a long-duration benchmark which comprises 10 long-horizon compositions, each formed by sequencing two LIBERO-90 goals, requiring multi-stage execution.
@@ C.8 Multimodality analysis
The Gumbel-Max trick ensures that token sampling still follows the logits’ softmax probability distribution, preserving diversity in the sampled results. To verify this, we conducted a qualitative evaluation of our model on the Push-T task Florence et al. (2021). In the initial stage, the control point, T-block, and target are placed symmetrically in fixed locations. MGP-Short then autoregressively generates 16 subsequent actions (4 tokens) and executes the first 8 actions based on the previous frame’s image and the control point position. The resulting trajectory plots (Fig. 9) show that MGP’s sampling still preserves the diversity distribution of the original data.